The Casings Plant of Wyoming Machines makes plastics shells for the companys calculators. (Each calculator requires one shell.) For each of the next two years, Wyoming expects to sell 660,000 calculators. The beginning finished goods inventory of shells at the Casings Plant is 90,000 units. However, the target ending finished goods inventory for each year is 25,000 units Each unit (shell) requires 6 ounces of plastic. At the beginning of the year, 250,000 ounces of plastic are in inventory. Management has set a target to have plastic on hand equal to two months sales requirements. Sales and production take place evenly throughout the year. Required: a. Compute the total targeted production of the finished product for the coming year n units b. Compute the required amount of plastic to be purchased for the coming year. (Do not round intermediate calculations.) to be p
Answer #1

a) Budgeted production in units = Sale unit+Desired ending inventory-Beginning inventory

= 660000+25000-90000

Budgeted production in units = 595000 Units

b) Material to be purchased in ounce = (595000*6)+(660000/12*2*6)-250000 = 3980000 Ounce